OpenStudy (anonymous):

When using u-substitution in integrating, we can't just introduce "new variables, right? so, like, integrating (sin x)^3, the u=sinx, and the du=cosx dx, and so if I were to use the u substitution, I would have to say (cosx)^-1 int(u^3)...would that work or no?

OpenStudy (anonymous):

$$ \int sin^3(x) $$ is a tough integration. Here is a video explaining it: https://www.youtube.com/watch?v=VeULk9R1z5E

OpenStudy (anonymous):

I had trouble with it when I started calc 2.

zepdrix (zepdrix):

Understand when you're making a substitution, you're replacing one variable with another. So yes, you can do something like u=sin x. You just can't replace your variable with the same variable: like x = sin x. Err maybe you could .. whatev.. that's not how we treat subs though XD Unfortunately u=sin x won't help in that particular problem. Oh oh we have an ODD power on sine, so this problem works out really nicely.

zepdrix (zepdrix):

\[\Large\rm \int\limits \sin^3x~dx=\int\limits (1-\cos^2x)\sin x~ dx\]Make the substitution u=cos x.

OpenStudy (anonymous):

^ you get this from the trig identity $$ sin^2(x) + cos^2(x) = 1 $$ $$ \therefore sin^2(x) = 1 - cos^2(x) $$

OpenStudy (anonymous):

OMAGOD THAT IS AMAZING

OpenStudy (anonymous):

but ok so introducing another x is NOT okay, right?

OpenStudy (anonymous):

You'll need to know your trig identities! Brush up if you are unfamiliar with them. You use them in ~half of Calc integrals.

OpenStudy (anonymous):

Well, I'm not sure if it's "legal" or not in this sense, but it's definitely not helpful. You don't want to introduce variables if they only make the integral harder.

zepdrix (zepdrix):

Yah don't introduce another x :U silly

zepdrix (zepdrix):

These integrals can be very easy or quite burdensome depending on what the integral looks like. For example:\[\Large\rm \int\limits \sin^{51}x~dx\]Would be MUCH easier to solve than,\[\Large\rm \int\limits \sin^6x~dx\]Even powers require a bit of fancy footwork.

OpenStudy (anonymous):

0-0 WATTTT a power of 51 is easier????

zepdrix (zepdrix):

hehe yah XD it's a simple u-sub. The power of 6 .. you need to either apply the Half-Angle Identity for Sine like ... 5 times, Or use the Sine Reduction Formula 3 times.

OpenStudy (anonymous):

To finish the integral, you would let u = cos(x). Therefore, du = -sin(x)dx -> which can be rewritten as -> -du = sin(x)dx And now do: $$ \int (1- u^2) -du $$ Pull the negative out (it's just a constant) $$ -\int (1- u^2) du $$ Distribute the du $$ -\int du + u^2 du $$ Split into two integrals $$ -[\int du + \int u^2 du] $$ Solve the integrals in terms of u $$ -[u + \frac{u^3}{3}] $$ Plug cos back in $$ -(cos(x)) + \frac{(cos(x))^3}{3} $$ Which can also be rearranged as $$\frac{(cos(x))^3}{3} -(cos(x)) $$

OpenStudy (anonymous):

Just some motivation, I was failing Calc 2 at the beginning, I didn't get any of this. STUDY AND PRACTICE!!! I worked my butt off, ended up with an A, and I should get an A in Calc 3. It's trivial once you run through a bunch of examples. Practice, practice, practice, makes perfect :)

OpenStudy (anonymous):

0-0 woah. so then the sin^51 is basically the same except a whole bunch of ... (1−u2)?

OpenStudy (anonymous):

hmmm LOL I have the AP exam wednesday ( that's tomorrow) and I haven't gotten too in-depth with these... i'm not feeling too good right now...

OpenStudy (anonymous):

Should I do it all tomorrow...sigh I have a 4 chemistry reports to turn in the very same day, so might as well do that... thanks, though!!! Ya'll are SUPER AWESOME

zepdrix (zepdrix):

for the 51st power? Break it up into 50 and 1,\[\Large\rm \int\limits \sin^{50}x \sin x~dx\]Then rewrite your 50 powers using the square identity,\[\Large\rm \int (\sin^{2}x)^{25}\sin x~dx\]\[\Large\rm \int\limits (1-\cos^2x)^{25}\sin x~ dx\]Then again we do our u-sub! :) u=cos x

zepdrix (zepdrix):

Chemistry? :O oh boy.

OpenStudy (anonymous):

DANG that is a very powerful property...and yes...chemistry...nbfhleugwiahokdsnbvjhfidsajnkcvbhf I am SOOOO done with chemistry

zepdrix (zepdrix):

XD

OpenStudy (anonymous):

any tips, perhaps, for chemistry? I have a final tomorrow...

zepdrix (zepdrix):

Hmm naw :c Chem was a little tough for me. I like the equations where stuff turns into other stuff and you gotta balance it. Those feel more math'y. But overall, chem felt like too much memorization >.<

OpenStudy (anonymous):

hahahahah LOL stuff turns into other stuffz LOL so true~ well, thanks ^^ I'll try to keep that in mind...stuffz=>other schtuffz HAHAHA

zepdrix (zepdrix):

XD